OpenStudy (itsbribro):

1.) What is the quotient written in scientific notation? 3*10^20 over 4*10^5 A. 7.5*10^3 B. 0.75*10^4 C. 7.5*10^14 D. 7.5*10^15 2.) What is the product written in scientific notation? (4.5*10^-8) (6*10^-4) A. 2.7*10^32 B. 2.7*10^-13 C. 2.7*10^-12 D. 2.7*10^-11

OpenStudy (anonymous):

3*10^20 over 4*10^5 is written as 7.5*10^14

OpenStudy (anonymous):

and... (4.5*10^-8) (6*10^-4) is written as 2.7*10^32. Took this lesson last week on math, k12.

OpenStudy (mayankdevnani):

\[\large \bf \frac{3 \times 10^{20}}{4 \times 10^5}\] Remember :- \[\large \bf \frac{10^a}{10^b}=10^{a-b}\] so,we get \[\large \bf \frac{3}{4} \times 10^{20-5}=0.75 \times 10^{15}\] then multiply and divide by 10, \[\large \bf 0.75 \times 10^{15} \times \frac{10}{10}=7.5 \times 10^{14}=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

so option C is correct :)

OpenStudy (mayankdevnani):

understood? @itsbribro

OpenStudy (itsbribro):

ohh ok thank you

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (mayankdevnani):

so can you do your last question self ? @itsbribro

OpenStudy (mayankdevnani):

*yourself

OpenStudy (itsbribro):

yes thank you

OpenStudy (mayankdevnani):

welcome